OpenStudy (anonymous):

for the given functions f and g, find the following. For parts (a)-(d), also find the domain. x=2x^2+3;g(x)=4x^3+1 for a (f+g)(x) and d (f/g)(x)

OpenStudy (anonymous):

\[(f+g)(x)=f(x)+g(x)\] so our job is to put a plus sign between them and then combine like terms you will get a polynomial, and domain will be all real numbers

OpenStudy (anonymous):

similarly \[\frac{f}{g}(x)=\frac{f(x)}{g(x)}\] so put f on top of g you will get \[\frac{f}{g}(x)=\frac{2x^2+3}{4x^3+1}\]

OpenStudy (anonymous):

for this one the denominator cannot be zero, so to find the domain you must solve \[4x^3+1\neq 0\] \[4x^3\neq -1\] \[x^3\neq -\frac{1}{4}\] \[x\neq -\frac{1}{\sqrt[3]{4}}\]

OpenStudy (anonymous):

ty